The perimeter of a rectangle is 25cm. The ratio of its side lengths is 2 to 3. What is the length of the shorter side of the rectangle, in cm?
* I do not need you to answer this! I just need a formula to solve it or step by step equation similar to this one.

Answer by richwmiller(17219)

2L+2W=P
2*(L+W)=P
2*(2x+3x)=25
2*(5x)=25
10x=25
x=2.5
2x=5
3x=7.5
the smaller side is 5 cm